# Hilbert’s Theorem 90, periodicity, and roots of Artin-Schreier polynomials

### v1

@@ Abstract
Let $E/F$ be a cyclic field extension of degree $n$, and let $\sigma$ generate the group $\textup{Gal}(E/F)$. If $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}\sigma^{i}y=0$, then the additive form of Hilbert’s Theorem 90 asserts that $y=\sigma x-x$ for some $x\in E$. Suppose that $E$ has characteristic $p$. We prove that $x$ gives rise to a periodic sequence $x_{0},x_{1},\dots$ which has period $pn_{p}$, where $n_{p}$ is the largest $p$-power that divides $n$. As an application, we find closed-form expressions for the roots of Artin-Schreier polynomials $t^{p}-t-y$. Let $y$ lie in the finite field ${\mathbb{F}}_{p^{n}}$ of order $p^{n}$. The Artin-Schreier polynomial $t^{p}-t-y\in{\mathbb{F}}_{p^{n}}[t]$ is reducible precisely when $\sum_{i=0}^{n-1}y^{p^{i}}=0$. In this case, $t^{p}-t-y=\prod_{k=0}^{p-1}(t-x-k)$ where $x=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ for some $z\in{\mathbb{F}}_{p^{e}}$ and $e=n_{p}$. The sequence $\left(\sum_{j=0}^{i-1}z^{p^{j}}\right)_{i\geqslant 0}$ is periodic with period $pe$, and if $e$ is small, then we give explicit $z$. Keywords: roots, reducible, Artin-Schreier polynomial, linearized polynomials 2020 Mathematics Subject Classification: 11T06, 11T55, 12E20
@@ 1 Introduction
This note began with the aims to show that there are infinitely many ‘quadratic formulas’ for quadratics over the finite field ${\mathbb{F}}_{2^{n}}$. The formulas have a periodicity, and the periodicity is related to the additive form of Hilbert’s Theorem 90 as we now explain.
Let $E/F$ be a cyclic field extension of degree $n$, and let $\sigma$ generate the Galois group $\textup{Gal}(E/F)$. If $\sum_{i=0}^{n-1}\sigma^{i}y=0$, then the additive form of Hilbert’s Theorem 90 asserts that $y=\sigma x-x$ for some $x\in E$. In this note, $E$ has prime characteristic $p$, and $n_{p}$ denotes the largest power of $p$ that divides $n$. We prove that $x$ can be chosen to have the form $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\sigma^{j}z\sigma^{i}y$ where $z$ is fixed by $\sigma^{n_{p}}$, and satisfies $\sum_{i=0}^{n-1}\sigma^{i}z=1$. Furthermore, $x=\sum_{i=0}^{n-1}x_{i}\sigma^{i}y$ where $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ defines a periodic sequence $x_{0},x_{1},\dots$ with period $pn_{p}$, see Theorem 3.1(b,c). A standard proof of the additive form of Hilbert’s Theorem 90 e.g. [4, p. 215] uses the ‘standard’ expression $x^{\prime}=R(z,y)$, see Lemma 3.2, where $y=x^{\prime}-\sigma(x^{\prime})$. This standard expression is very natural in one sense, see the preamble to Lemma 3.2. However, it makes questions about periodicity unnatural. Lemma 3.2 relates $x=R(y,z)$ to $x^{\prime}=R(z,y)$ via the equation $x+x^{\prime}=-\sum_{i=0}^{n-1}\sigma^{i}(yz)$.
@@
In some cases the generator $\sigma\colon E\to E$ may be taken to be the $p$th power map $\sigma(x)=x^{p}$, e.g. if $E={\mathbb{F}}_{p^{n}}$ is the finite field of order $p^{n}$ and $F={\mathbb{F}}_{p}$. If $\sum_{i=0}^{n-1}y^{p^{i}}=0$ and $\sum_{i=0}^{n-1}z^{p^{i}}=1$, then $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ is a root of the Artin-Schreier polynomial $t^{p}-t-y$. For example, if $n_{p}=1$, we may take $z=\frac{1}{n}$ and $x_{i}=\frac{i}{n}$ defines a sequence of period $p$. Furthermore, $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{p^{i}}$ is a closed-form expression for a root of $t^{p}-t-y$ in $E$ when $\sum_{i=0}^{n-1}y^{p^{i}}=0$. The case $n_{p}=p$ uses a different choice for $z$, and $n_{p}=p^{2}$ a different choice, etc. This suggests that there are infinitely many formulas for the roots of the polynomial $t^{p}-t-y$ over the finite field ${\mathbb{F}}_{p^{n}}$ that depend on the $p$-part $n_{p}$ of $n$. Choosing $z$ in the field ${\mathbb{F}}_{p^{e}}$ with $e=n_{p}$ and $\sum_{i=0}^{n-1}z^{p^{i}}=1$ is easy in practise: with probability $1-\frac{1}{p}$ a random element $\zeta\in{\mathbb{F}}_{p^{e}}$ has $\tau\coloneq\sum_{i=0}^{n-1}\zeta^{p^{i}}\neq 0$, so take $z=\tau^{-1}\zeta$. If $n_{p}$ is fixed, and we know $z\in{\mathbb{F}}_{p^{e}}$ satisfying $\sum_{i=0}^{n-1}z^{p^{i}}=1$, then $R(y,z)$ may be viewed as a closed-form expression for a root of $t^{p}-t-y$. For examples with $p=2$ see Lemma 4.8.
@@ 2 Background
This work fits into a very classical setting, namely generalizing Galois theory to characteristic $p$. The following important result follows from the additive form of Hilbert’s Theorem 90, see [2, Theorem 4.33] and [4, VIII, Theorem 11].
@@
Let $F$ be field of prime characteristic $p$.
@@
    Given $y\in F$, the polynomial $t^{p}-t-y$ is either irreducible over $F$, or it has $p$ distinct roots $x+a$ where $a\in{\mathbb{F}}_{p}$ and $x\in F$.
The trace map $\textup{Tr}_{F}^{E}(x)=\sum_{\alpha\in\textup{Gal}(E/F)}\alpha(x)$ of a finite Galois extension $E/F$, is an $F$-linear map $E\to F$ which is nonzero by Artin’s Lemma [4, VIII, Theorem 7]. In certain cases, the trace map can be used to distinguish whether or not $t^{p}-t+y$ is irreducible.
@@
If $\textup{Tr}_{F}^{E}(y)=0$, then by Hilbert’s Theorem 90, $y=z-\sigma(z)=z-z^{p}$ for some $z\in E$. Hence $t^{p}-t+y$ is reducible over $E$. Conversely, suppose if $t^{p}-t+y$ is reducible over $E$, then $t^{p}-t+y$ has a root $z\in E$, by Theorem 2.1(b). Using $\sigma z=z^{p}$ shows
@@
Finally, set $y=1$. Then $\textup{Tr}_{F}^{E}(1)=p\cdot 1=0$ and some $\zeta\in E$ satisfies $t^{p}-t+1=0$. However, $\sigma\zeta=\zeta^{p}=\zeta-1$ implies $\sigma^{i}\zeta=\zeta^{p}=\zeta-i$ so $\zeta$ has $p$ distinct conjugates in $E$ and so $E=F(\zeta)$. Therefore $E=F(z)$ where $z=\zeta^{-1}$ is a root of the (reverse) polynomial $t^{p}-t^{p-1}+1$. Since the conjugates $z,\sigma z,\dots,\sigma^{p-1}z$ are pairwise distinct, $t^{p}-t^{p-1}+1$ is also irreducible over $F$. The coefficient of $t^{p-1}$ shows $\textup{Tr}^{E}_{F}(z)=1$, as claimed. ∎
@@
[…]
If $E={\mathbb{F}}_{2^{n}}$ is a finite field of order $2^{n}$, then $x^{2}+bx+c$ has a double root $x=c^{2^{n-1}}$ if $b=0$ because $c^{2^{n}}=c$ holds. If $b\neq 0$, then $x^{2}+bx+c=b^{-2}\left(t^{2}+t+b^{-2}c\right)$ where $t=b^{-1}x$. Taking $y=b^{-2}c$, the quadratic $t^{2}+t+y$ has roots $R(y,z),R(y,z)+1$ where $R(y,z)$ is given by the periodic formula $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{2^{j}}y^{2^{i}}$, see Lemma 4.8.
@@ 3 Periodicity and Hilbert’s additive Theorem 90
If $E/F$ is a finite Galois extension and $\sigma\in\textup{Gal}(E/F)$, then $\textup{Tr}^{E}_{F}(\sigma x)=\textup{Tr}^{E}_{F}(x)$ holds. Hence if $y\in E$ has the form $\sigma x-x$ for some $x\in E$ and $\sigma\in\textup{Gal}(E/F)$, then $\textup{Tr}^{E}_{F}(y)=0$ holds. If $\textup{Gal}(E/F)$ is finite and cyclic, then the converse holds by Hilbert’s Theorem 90. Throughout this section $E/F$ is a cyclic Galois extension of degree $n$ and prime characteristic $p$, and $\textup{Gal}(E/F)$ is generated by $\sigma$. Hence $\sigma^{n}=1$ and $\textup{Tr}_{F}^{E}(y)=\sum_{i=0}^{n-1}\sigma^{i}y$.
We prove in Theorem 3.1 that the element $x$ has an intrinsic periodicity. This observation is used in Corollary 4.1 to give a polynomial formula for the roots of (reducible) Artin-Schreier polynomials.
Theorem 3.1.
  1. (a)
    $y=\sigma x-x$ where $x=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$ for some $z\in E$ with $\textup{Tr}^{E}_{F}(z)=1$;
  2. (b)
  3. (c)
Proof.
(a) Suppose that $x$ has the form $\sum_{i=0}^{n-1}x_{i}\sigma^{i}y$ where $x_{0},\dots,x_{n-1}\in E$. Then $\sigma x-x=-x_{0}y+\sum_{i=1}^{n-1}(\sigma x_{i-1}-x_{i})\sigma^{i}y+\sigma x_{n-1}\sigma^{n}y$. Assume for the moment that there exists an element $z\in E$ satisfying
@@
We prove that these necessary equations are sufficient. First, the trace map $\textup{Tr}_{F}^{E}\colon E\to F$ is surjective, so we may choose $z\in E$ so that $\textup{Tr}_{F}^{E}(z)=1$. Now define $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where the empty sum $x_{0}=\sum_{j=1}^{0}\sigma^{j}z$ is 00 by convention. Then the $n$ equations in (1) hold, and $\sigma x-x$ equals $y$.
(b) First, we show that the fixed subfield $K$ of $\sigma^{n_{p}}$ contains an element $z$ satisfying $\textup{Tr}^{E}_{F}(z)=1$. Since $n_{p^{\prime}}\neq 0$ in $F$ and the map $\textup{Tr}^{K}_{F}\colon K\to F$ is surjective, there exists an element $z\in K$ such that $\textup{Tr}^{K}_{F}(z)=(n_{p^{\prime}})^{-1}$. Then $\textup{Tr}^{E}_{K}(z)=|E:K|z=n_{p^{\prime}}z$ and hence $\textup{Tr}^{E}_{F}(z)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(z))=n_{p^{\prime}}\textup{Tr}^{K}_{F}(z)=1$. Next, suppose that $L=F(z)$ and $\textup{Tr}^{E}_{F}(z)=1$. Then $1=\textup{Tr}^{L}_{F}(\textup{Tr}^{E}_{L}(z))=|E:L|\textup{Tr}^{L}_{F}(z)$. Therefore $p\nmid|E:L|$, so $n_{p}=|K:F|$ divides $e=|L:F|$ as claimed. In summary, the minimal polynomial over $F$ of an element $z$ with $\textup{Tr}^{E}_{F}(z)=1$ has degree a multiple of $n_{p}$, and degree precisely $n_{p}$ if $z$ lies in $K\setminus K_{0}$ where $K_{0}$ is the unique (maximal) subfield of $K$ with $F\leqslant K_{0}<K$ and $|K:K_{0}|=p$.
(c) Let $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where $\textup{Tr}^{E}_{F}(z)=1$. Let $L=F(z)$ and $e=|L:F|$. Hence $x_{e}=\textup{Tr}^{L}_{F}(z)\neq 0$ by part (b) and $\sigma^{j}z=z$ implies $e\mid j$. Suppose that $x_{k}=0$ where $0\leqslant k<pe$ and write $k=ie+j$ where $0\leqslant i<p$ and $0\leqslant j<e$. The equation $x_{ie+j}=ix_{e}+x_{j}$ shows that $x_{0}=x_{pe}=0$. Thus $x_{k}=ix_{e}+x_{j}=0$ and $x_{e}\in F$ gives
$x_{j+1}-z=\sigma x_{j}=\sigma(-ix_{e})=-ix_{e}=x_{j}.$
The attentive reader will notice that the proof of Theorem 3.1(a) differs from the standard proof: a symmetry has been broken as we now explain. There are two versions of theorems regarding the kernels of the norm and trace maps. The former is usually called Hilbert’s (multiplicative) Theorem 90 because of the numbering in [1, p. 272], and the latter the additive form of Hilbert’s Theorem 90. Emmy Noether generalized the method used to prove these theorems, essentially by introducing 1-cocyles, to handle a noncyclic (finite) Galois group $G\coloneq\textup{Gal}(E/F)$, see [2, Theorems 4.29 and 4.30]. Her first result [2, Theorem 4.29] says: given a function $y\colon G\to E\setminus\{0\}$ satisfying $y_{\sigma\tau}=y_{\sigma}\sigma(y_{\tau})$ for $\sigma,\tau\in G$, there exists an $x\in E\setminus\{0\}$ such that $y_{\tau}=x\tau(x)^{-1}$. Her second result [2, Theorem 4.30] says: given a function $y\colon G\to E$ satisfying $y_{\sigma\tau}=\sigma(y_{\tau})+y_{\sigma}$ for $\sigma,\tau\in G$, there exists an $x\in E$ such that $y_{\tau}=x-\tau(x)$. When $G=\langle\sigma\rangle$ is cyclic order $n$, the 1-cocycle conditions suggest the definitions $y_{\phi^{i}}=y(\sigma y)\cdots(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{1}=y$, and similarly in the additive case $y_{\phi^{i}}=y+(\sigma y)+\cdots+(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{0}=y$. This gives rise to the (multiplicative and additive) formulas $x=\sum_{i=0}^{n-1}\left(\prod_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$ and $x=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$, respectively. The additive expression is $R(z,y)$ where $R(y,z)$ is in Theorem 3.1(a). The expressions $R(z,y)$ and $R(y,z)$ are related as follows.
@@
The equations $\sum_{j=0}^{n-1}\sigma^{j}z=1$ and $\sum_{i=0}^{n-1}\sigma^{i}y=0$ imply
@@
Swapping the order of summation, swapping $i\leftrightarrow j$, and noting $\textup{Tr}^{E}_{F}(zy)=\sum_{i=0}^{n-1}\sigma(zy)$ gives

### v2

@@ Abstract
Let $E/F$ be a cyclic field extension of degree $n$, and let $\sigma$ generate the group $\textup{Gal}(E/F)$. If $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}\sigma^{i}y=0$, then the additive form of Hilbert’s Theorem 90 asserts that $y=\sigma x-x$ for some $x\in E$. When $E$ has characteristic $p>0$ we prove that $x$ gives rise to a periodic sequence $x_{0},x_{1},\dots$ which has period $pn_{p}$, where $n_{p}$ is the largest $p$-power that divides $n$. We also show, if $y$ lies in the finite field ${\mathbb{F}}_{p^{n}}$, then the roots of a reducible Artin-Schreier polynomial $t^{p}-t-y$ have the form $x+u$ where $u\in{\mathbb{F}}_{p}$ and $x=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ for some $z\in{\mathbb{F}}_{p^{e}}$ with $e=n_{p}$. Furthermore, the sequence $\left(\sum_{j=0}^{i-1}z^{p^{j}}\right)_{i\geqslant 0}$ is periodic with period $pe$. Keywords: Hilbert’s Theorem 90, Artin-Schreier polynomials, periodic sequence 2020 Mathematics Subject Classification: 11T06, 12F10, 11T55
@@ 1 Introduction
Given a quadratic equation $ax^{2}+bx+c=0$ over a finite field ${\mathbb{F}}_{2^{n}}$ of order $2^{n}$, Chen [1] gives ‘formulas’ for the roots that involve $a,b,c$ and some element of trace 1. We give a unified treatment of these formulas which in prime characteristic $p$ depends on $n_{p}$, the largest $p$-power divisor of $n$. This is related the additive version of Hilbert’s Theorem 90 and to sequences of period $pn_{p}$. We say that a sequence $(x_{i})_{i\geqslant 0}$ has period $e$ if $x_{i}=x_{j}$ whenever $i\equiv j\pmod{e}$, and $e$ is minimal. In particular, we prove that a root of a reducible quadratic $x^{2}+x+y\in{\mathbb{F}}_{2^{n}}[x]$ has the form $x=\sum_{i=0}^{n-1}x_{i}y^{2^{i}}$, $x_{i}\in{\mathbb{F}}_{2^{n}}$, where the sequence $(x_{i})_{i\geqslant 0}$ has period $2n_{2}$. If $n_{2}=1$, we may take $x_{i}=0,1$ whenever $i\equiv 0,1\pmod{2}$, and if $n_{2}=2$, we may take $x_{i}=0,\omega,1,\omega^{2}$ whenever $i\equiv 0,1,2,3\pmod{4}$: here $\omega^{2}+\omega+1=0$ and $\omega$ is a primitive cube root of 1. The periodic sequence $(x_{i})_{i\geqslant 0}$ is described in detail, and for larger values of $n_{2}$, in Lemma 4.7.
This note proves the following two main results which we contextualize below. Given a finite Galois extension $E/F$ with group $\Gamma$, the trace map $E\to F$ is $\textup{Tr}^{E}_{F}(y)=\sum_{\gamma\in\Gamma}\gamma y$.
Theorem 1.1.
Let $E/F$ be a cyclic extension of degree $n$. Suppose $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $y,z\in E$ satisfy $\textup{Tr}^{E}_{F}(y)=0$, $\textup{Tr}^{E}_{F}(z)=1$, then $y=\sigma x-x$ where $x=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$.
The element $z$ in Theorem 1.1 satisfies additional properties in prime characteristic.
Theorem 1.2.
  1. (a)
  2. (b)
@@
Polynomials of the form $t^{p}-t-y$ where $y\in E$ were studied by Artin and Schreier in the context of generalizing Galois theory to characteristic $p$. Our theorems relate to roots of these Artin-Schreier polynomials. Suppose $\textup{char}(E)=p>0$ and the $p$th power map $\sigma\colon x\mapsto x^{p}$ of $E$ generates $\textup{Gal}(E/F)$, e.g. when $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$. If $\textup{Tr}^{E}_{F}(y)=0$ and $\textup{Tr}^{E}_{F}(z)=1$, then $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ is a root of the polynomial $t^{p}-t-y$. For example, if $n_{p}=1$, we may take $z=\frac{1}{n}$ and $x_{i}\coloneq\sum_{j=0}^{i-1}z^{p^{j}}=\frac{i}{n}$ defines a sequence of period $p$, and $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{p^{i}}$ is a root of $t^{p}-t-y$. Because different choices of $n_{p}$ involve different choices for $z$ with $\textup{Tr}^{E}_{F}(z)=1$, this suggests that there are infinitely many (polynomial) formulas for the roots of polynomials $t^{p}-t-y\in E[t]$. If $E={\mathbb{F}}_{p^{n}}$, $F={\mathbb{F}}_{p}$ and $e=n_{p}$, then finding $z$ in ${\mathbb{F}}_{p^{e}}$ with $\textup{Tr}^{E}_{F}(z)=1$ is easy in practise: with probability $1-\frac{1}{p}$ a random element $\zeta\in{\mathbb{F}}_{p^{e}}$ has $\tau\coloneq\textup{Tr}^{E}_{F}(\zeta)\neq 0$, so $z=\tau^{-1}\zeta$ has $\textup{Tr}^{E}_{F}(z)=1$.
There are two versions of Hilbert’s Theorem 90: the additive form and the multiplicative form: we abbreviate these AH90 and MH90, respectively. Theorem 1.1 is different from usual version of AH90 in that the formula for $x$ is very slightly different. This difference leads to Theorem 1.2. We remark first that MH90 is so numbered because it appears as Satz 90 in [2] on p. 272, and Hilbert never actually proved AH90. The standard proof of AH90 is given in Lang [5, p. 215] and it follows Emil Artin’s ideas. In Theorem 1.1, $x$ equals $R(y,z)\coloneq\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\sigma^{j}z\sigma^{i}y$ where $z\in E$ satisfies $\textup{Tr}^{E}_{F}(z)=1$ and $y=\sigma x-x$ holds. By contrast, the standard proof of AH90 uses a different element, namely $x^{\prime}=R(z,y)$ and $y=x^{\prime}-\sigma x^{\prime}$ holds. The proofs of AH90 I have found, assume $\textup{Tr}^{E}_{F}(z)\neq 0$ (but renormalizing to get $\textup{Tr}^{E}_{F}(z)=1$ is easy) and show that
$x^{\prime}$ $=\frac{1}{\textup{Tr}^{E}_{F}(z)}\left(y\sigma(z)+(y+\sigma(y))\sigma^{2}(z)+\cdots+(y+\sigma(y)+\cdots+\sigma^{n-2}(y))\sigma^{n-1}(z)\right)$
$=\frac{1}{\textup{Tr}^{E}_{F}(z)}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\sigma^{j}y\sigma^{i}z=\frac{1}{\textup{Tr}^{E}_{F}(z)}R(z,y)$
satisfies $y=x^{\prime}-\sigma x^{\prime}$. The point here is that writing $x$ in the form $\sum_{i=0}^{n-1}x_{i}\sigma^{i}y$ for some $x_{i}\in E$, suggests studying the sequence $(x_{i})_{i\geqslant 0}$, whereas the form of $x^{\prime}$ does not.
We remark that the expression $x^{\prime}=R(z,y)$ is very natural in the context of Galois cohomology. Indeed, this form suggests itself via Emmy Noether’s equations, and these equations lead to analogous proofs of AH90 and MH90 as we explain. Emmy Noether generalized the method used to prove AH90 and MH90 by introducing 1-cocyles, to handle a noncyclic (finite) Galois group $\Gamma\coloneq\textup{Gal}(E/F)$, see [3, Theorems 4.29 and 4.30]. Her result [3, Theorem 4.29] says: given a function $y\colon\Gamma\to E\setminus\{0\}$ satisfying $y_{\beta\gamma}=y_{\beta}\beta(y_{\gamma})$ for $\beta,\gamma\in\Gamma$, then for some $z\in E$, $x=\sum_{\beta\in\Gamma}y_{\beta}\beta(z)$ satisfies $y_{\gamma}=x\gamma(x)^{-1}$ for all $\gamma\in\Gamma$. Her result [3, Theorem 4.30] says: given a function $y\colon\Gamma\to E$ satisfying $y_{\beta\gamma}=y_{\beta}+\beta(y_{\gamma})$ for $\beta,\gamma\in\Gamma$, then for some $z\in E$, $x=\textup{Tr}^{E}_{F}(z)^{-1}\sum_{\beta\in\Gamma}y_{\beta}\beta(z)$ satisfies $y_{\gamma}=x-\gamma(x)$ for all $\gamma\in\Gamma$. When $\Gamma=\langle\sigma\rangle$ is cyclic of order $n$, the 1-cocycle conditions imply that $y_{\sigma^{i}}=y(\sigma y)\cdots(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{1}=y$, and similarly in the additive case $y_{\sigma^{i}}=y+(\sigma y)+\cdots+(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{1}=y$. This gives rise to the multiplicative formula $x=\sum_{i=0}^{n-1}\left(\prod_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$ and when $\textup{Tr}^{E}_{F}(z)=1$, the additive formula $x=R(z,y)\coloneq\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$, respectively. Theorem 1.1 uses the related expression $R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$. Lemma 3.1 in some sense measures the asymmetry between $R(y,z)$ and $R(z,y)$ via the equation $R(y,z)+R(z,y)=-\sum_{i=0}^{n-1}\sigma^{i}(yz)$.
@@ 2 Background
This work fits into a very classical setting, namely generalizing Galois theory to characteristic $p$. The following important result follows from the AH90, see [3, Theorem 4.33] and [5, VIII, Theorem 11] for proofs of Theorem 2.1.
@@
Let $E$ be field of prime characteristic $p$.
@@
    Given $y\in E$, the polynomial $t^{p}-t-y$ is either irreducible over $E$, or it has $p$ distinct roots $x+u$ where $x\in E$ and $u\in{\mathbb{F}}_{p}$.
The trace map $\textup{Tr}_{F}^{E}(x)=\sum_{\gamma\in\textup{Gal}(E/F)}\gamma(x)$ of a finite Galois extension $E/F$, is an $F$-linear map $E\to F$ which is nonzero by Artin’s Lemma [5, VIII, Theorem 7]. In certain cases, the trace map can be used to distinguish whether or not $t^{p}-t+y$ is irreducible.
@@
If $\textup{Tr}_{F}^{E}(y)=0$, then $y=z-\sigma(z)=z-z^{p}$ for some $z\in E$ by AH90. Hence $t^{p}-t+y$ is reducible over $E$. Conversely, suppose if $t^{p}-t+y$ is reducible over $E$, then $t^{p}-t+y$ has a root $z\in E$, by Theorem 2.1(b). Using $\sigma z=z^{p}$ shows
@@
Finally, set $y=1$. Then $\textup{Tr}_{F}^{E}(1)=p\cdot 1=0$ and some $\zeta\in E$ satisfies $t^{p}-t+1=0$. However, $\sigma\zeta=\zeta^{p}=\zeta-1$ implies $\sigma^{i}\zeta=\zeta-i$ so $\zeta$ has $p$ distinct conjugates in $E$ and $E=F(\zeta)$. Therefore $E=F(z)$ where $z=\zeta^{-1}$ is a root of the (reverse) polynomial $t^{p}(t^{-p}-t^{-1}+1)=t^{p}-t^{p-1}+1$. Since the conjugates $z,\sigma z,\dots,\sigma^{p-1}z$ are pairwise distinct, $t^{p}-t^{p-1}+1$ is also irreducible over $F$. The coefficient of $t^{p-1}$ shows that $\textup{Tr}^{E}_{F}(z)=1$. ∎
We note that only part of Theorem 2.1(b) generalizes. If ${\mathbb{F}}_{q}\leqslant F\leqslant E$, and $t^{q}-t-y$ has a root $x\in E$, then $t^{q}-t-y$ factors as $\prod_{a\in{\mathbb{F}}_{q}}(t-(x+a))$. However, if $t^{q}-t-y$ has no root in $E$, we can not conclude that $t^{q}-t-y$ is irreducible over $E$. For example, if $q=4$ then $t^{4}+t+1$ factors as $(t^{2}+t+\omega)(t^{2}+t+\omega^{2})$ over ${\mathbb{F}}_{4}={\mathbb{F}}_{2}(\omega)$.
@@ 3 Periodicity and Hilbert’s additive Theorem 90
Let $E/F$ be a Galois extension with group $\Gamma=\textup{Gal}(E/F)$. Thus $\textup{Tr}^{E}_{F}(\sigma x)=\textup{Tr}^{E}_{F}(x)$ for $\sigma\in\Gamma$. Hence if $y\in E$ has the form $\sigma x-x$ for some $x\in E$ and $\sigma\in\Gamma$, then $\textup{Tr}^{E}_{F}(y)=0$. If $\Gamma$ is finite and cyclic, then the converse holds by AH90. In this section $E/F$ is a cyclic extension of degree $n$, and $\textup{Gal}(E/F)=\langle\sigma\rangle$. Thus $\sigma^{n}=1$ and $\textup{Tr}_{F}^{E}(y)=\sum_{i=0}^{n-1}\sigma^{i}y$.
Proof of Theorem 1.1.
Suppose that $x$ has the form $\sum_{i=0}^{n-1}x_{i}\sigma^{i}y$ where $x_{0},\dots,x_{n-1}\in E$. We aim to solve for the coefficients $x_{i}$. Observe that
$\sigma x-x=-x_{0}y+\sum_{i=1}^{n-1}(\sigma x_{i-1}-x_{i})\sigma^{i}y+\sigma x_{n-1}\sigma^{n}y.$
Assume for the moment that there exists an element $z\in E$ satisfying
@@
To see that these necessary equations are sufficient, choose $z\in E$ so that $\textup{Tr}_{F}^{E}(z)=1$ which is possible as the trace map $\textup{Tr}_{F}^{E}\colon E\to F$ is surjective. Next, define $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where the empty sum $x_{0}=\sum_{j=1}^{0}\sigma^{j}z$ is 00 by convention. Then the $n$ equations in (1) hold, and $\sigma x-x$ equals $y$, as claimed. ∎
@@
The equations $\sum_{i=0}^{n-1}\sigma^{i}y=0$ and $\sum_{j=0}^{n-1}\sigma^{j}z=1$ imply
@@
Swapping the order of summation, swapping $i\leftrightarrow j$, and noting $\textup{Tr}^{E}_{F}(zy)=\sum_{i=0}^{n-1}\sigma^{i}(zy)$ gives